OpenStudy (anonymous):

x2 + y2- 4x + 2y - 11 = 0 Classify the conic section and write its equation in standard form.

OpenStudy (anonymous):

this is equation of a circle x^2 - 4x + y^2 + 2y - 11 = 0 (x - 2)^2 + (y + 1)^2 - 4 - 1 - 11 = 0 standard form : (x - 2)^2 + ( y + 1)^2 = 16

OpenStudy (anonymous):

do you need any more help?